OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x

OpenStudy (anonymous):

3/4x+5/8=4x subtract 3/4x from both sides 5/8=3.25x divide by 3.25 x=5/26

OpenStudy (anonymous):

:)